OpenStudy (anonymous):

On a map of the United States, one inch represents 400 miles. If New York and California are six and 3 4 inches apart, how many miles apart are they?

OpenStudy (anonymous):

Welcome to OpenStudy! :) \(\sf\large 6 \frac{3}{4}=\frac{27}{4}\) So now just multiply \(\sf\frac{27}{4}\times 400=?\)

OpenStudy (anonymous):

2700, got it. thank you so much

OpenStudy (anonymous):

Anytime :)

TheSmartOne (thesmartone):

Good job @iamrobin ;) @karyymartinez 2700 is correct and \(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0